OpenStudy (anonymous):

log x + 5^36 = 2 x = 1 x = 11 x = 13 x = 23

OpenStudy (anonymous):

I think that there is either something missing or I'm misinterpreting something. I'm assuming that the log in use here is base 10. In this case: \[\log x=2-5^{36}\]Then\[10^{2-5^{36}} = x = 10^{2}/10^{5^{36}}\]It is fairly easy to see that x is arbitrarily close to zero, leading me to believe that something is awry here.